# Is Doublet-Triplet Splitting Necessary?

### v3

@@ Abstract
We demonstrate that it is not necessary to substantially break the mass degeneracy between the Standard Model Higgs boson doublet and the corresponding scalar leptoquark, where these two fields comprise a single five-dimensional $SU(5)$ representation. More precisely, we show that the experimental data on partial proton decay lifetimes cannot place any meaningful bound on the mass of the scalar leptoquark, a.k.a. the color triplet, if one includes effects of higher-dimensional operators within the Georgi-Glashow model. We also point out that the complete suppression of proton decay via the scalar leptoquark mediation is compatible with the partial suppression of the gauge mediated two-body proton decay signatures. Finally, we derive a proper limit on the cutoff scale of the aforementioned higher-dimensional operators and outline how this scenario could be tested.
@@ I Introduction
One of the most prominent features of the the Georgi-Glashow model Georgi:1974sy is an ever-present need to introduce an enormous splitting between the masses of the two multiplets residing in the same five-dimensional scalar representation $5_{H}$ of the theory. One of these two multiplets is the Standard Model (SM) Higgs boson doublet $D$, whereas the other one is the color triplet $T$ that generates proton decay. Since the Higgs boson is light while the proton decay mediating scalar leptoquark should be extremely heavy, one needs to strongly break mass degeneracy between them. This issue has been referred to in the literature as the ($D$)oublet-($T$)riplet splitting problem.
@@
We demonstrate that it is not necessary to accomplish the aforementioned mass split if one includes effects of the higher-dimensional operators within the Georgi-Glashow model. The inclusion of these operators allows one to completely suppress proton decay signatures via the scalar leptoquark mediation without affecting viability of the mechanism that generates masses of the SM charged fermions. We furthermore show that this intriguing possibility is also compatible with the partial suppression of the gauge mediated two-body proton decay signatures.
@@
We introduce the most relevant details of the $SU(5)$ setup and the accompanying notation in Sec. II. The proton decay suppression mechanisms for both the scalar and gauge mediations are discussed in detail in Sec. III. We subsequently show how to establish an accurate upper bound on the cutoff scale associated with the higher-dimensional operators if one resorts to these suppression mechanisms in Sec. IV. Final remarks are presented in Sec. V and all our findings summarized in Sec. VI.
@@ II Setup
$+\frac{1}{\Lambda}Y_{3}^{\alpha\beta}24_{Hm}^{n}\epsilon_{ijkln}+\frac{1}{\Lambda}Y_{4}^{\alpha\beta}24_{Hi}^{n}\epsilon_{jklmn}\bigg{\}}+\mathrm{h.c.},$ (1)
@@
The vacuum expectation values (VEVs) of relevance are
@@
Mass matrices of the SM fermions are diagonalized via
@@ III Proton Decay Suppression
Since the linear combinations of Yukawa coupling matrices that enter the charged fermion masses $M_{E}$, $M_{D}$, and $M_{U}$ differ from those that are featured in the interactions of the scalar leptoquark $T$ with the SM fermions, one can suppress the latter without affecting viability of the former. For example, one can set to zero all quark-quark couplings of $T$, as given in Eqs. (13) and (15), by imposing the following three relations
$Y_{2}\epsilon$ $=Y_{1}\epsilon-Y_{d}$ (17)
@@
where the viability of Eq. (18) will be discussed in detail later on. These assumptions consequentially lead to
@@
Even though we still need to demonstrate that one can indeed cancel $d=4$ contributions with $d=5$ contributions in the up-type quark sector with perturbative Yukawa couplings, Eqs. (17) through (23) already imply that generation of viable charged fermion masses is compatible with complete suppression of scalar mediated proton decay signatures. Therefore, one of the main results of this manuscript is that experimental data on partial proton decay lifetimes cannot provide any meaningful constraint on the scalar leptoquark mass if one includes effects of higher-dimensional operators. Our result is applicable to any non-supersymmetric $SU(5)$ setup that ($i$) uses a single $5_{H}$ and ($ii$) resorts to $d=5$ operators to provide viable masses for the SM charged fermions.
@@
Although proton decay via gauge boson exchange cannot be completely rotated away in $SU(5)$ Nandi:1982ew , it is possible to significantly suppress coefficients in Eqs. (27), (28), and (29Dorsner:2004xa . This, for example, can be achieved with the followings set of conditions Dorsner:2004xa
@@ IV Cutoff Scale Limit
For our purposes it is sufficient to take $M_{U}^{\mathrm{diag}}=\mathrm{diag}(0,0,m_{t})=\mathrm{diag}(0,0,y_{t}v_{5}/\sqrt{2})$, where $m_{t}$ and $y_{t}$ are the top quark mass and Yukawa coupling at $M_{\mathrm{GUT}}$ scale, respectively. This decomposition allows us to find a lower limit on the largest possible entry in $A$, i.e., $\max|(A)_{\alpha\beta}|$, in units of $y_{t}$. Since the elements of $A$ can only originate from the very last operator in Eq. (7), our approach should yield an accurate lower limit on parameter $\epsilon\equiv v_{24}/\Lambda$ if we insist on perturbativity of Yukawa couplings. Coincidentally, decomposition of Eq. (34) is compatible with our assumptions of Eqs. (18) and (19). This, in turn, will enable us to explicitly demonstrate that one can cancel $d=4$ contributions with $d=5$ contributions with perturbative couplings in the up-type quark sector.
@@
Here, $c_{\theta}\equiv\cos\theta$ and $s_{\theta}\equiv\sin\theta$, where $\theta$ is a free parameter, while $\eta_{1}$, $\eta_{2}$, $\eta_{3}$, and $\eta_{4}$ are four arbitrary phases. We can, furthermore, represent unitary matrix $U$ via
@@
To proceed, one would need to establish what the viable values of $v_{24}$ and $y_{t}$, to be used as input in Eq. (39), are. We accordingly observe that $v_{24}=M_{\mathrm{GUT}}\sqrt{(2\alpha_{\mathrm{GUT}}^{-1})/25\pi}$, where $\alpha_{\mathrm{GUT}}=g_{\mathrm{GUT}}^{2}/(4\pi)$. Also, a model Dorsner:2005fq ; Dorsner:2005ii ; Dorsner:2006hw that $(i)$ uses perturbative couplings to generate neutrino masses, $(ii)$ relies on operators of Eq. (1) to generate charged fermion masses, and $(iii)$ implements Eqs. (30), (31), and (32) to suppress gauge mediated proton decay signatures, yields $\alpha_{\mathrm{GUT}}^{-1}=37.1$ and $y_{t}=0.42$, where $\alpha_{\mathrm{GUT}}$ is evaluated via the two-loop level gauge coupling unification analysis, whereas $y_{t}$, at $M_{\mathrm{GUT}}$, is obtained through the one-loop level renormalisation group equation running. If these values for $\alpha_{\mathrm{GUT}}$ and $y_{t}$ are inserted into Eq. (39) we find that the upper limit on $\Lambda$ for a model of Ref. Dorsner:2005fq reads $\Lambda<57M_{\mathrm{GUT}}$, where we assume that $5/2\max|(Y_{4})_{\alpha\beta}|=\sqrt{4\pi}$. This bound on $\Lambda$, as inferred from Eq. (39), is another key finding of this manuscript. It is, as already advocated, much more stringent than the $\Lambda\lesssim 900M_{\mathrm{GUT}}$ bound reported in Ref. Dorsner:2006hw .
@@
Let us, as a qualitative example, assume that $\epsilon=0.05$. This, for $\alpha_{\mathrm{GUT}}^{-1}=37.1$ and $y_{t}=0.42$, implies that $\Lambda=19.4M_{\mathrm{GUT}}$, $2.25\max|(Y_{3}+Y_{3}^{T})_{\alpha\beta}|>$ 2.2, $2.45\max|(Y_{4})_{\alpha\beta}|/2>$ 2.4, and $0.882\max|(Y_{u}+Y_{u}^{T})_{\alpha\beta}|>$ 0.88 if one is to implement suppression of both scalar and gauge mediations of proton decay.
@@
V Final Remarks
Final comments are in order.
  • Our derivation of the limit on $\Lambda$ is accurate up to the corrections of the order of $m_{c}/m_{t}$, where $m_{c}$ is the charm quark mass. Moreover, the values of elements in $V_{\mathrm{CKM}}$ used to generate bounds in Eqs. (39), (40), and (41) are sourced from the Particle Data Group ParticleDataGroup:2022pth , although the only $V_{\mathrm{CKM}}$ entry that matters for our analysis is the Cabibbo angle.
  • We have explicitly demonstrated how to suppress the quark-quark couplings of the scalar leptoquark $T$ in Sec. III. We could have alternatively opted to suppress the quark-lepton couplings of Eqs. (12), (14), and (16) via $Y_{2}\epsilon=\frac{2}{3}\left(Y_{d}-Y_{1}\epsilon\right)$ and $\left(Y_{3}+Y_{3}^{T}\right)\epsilon=\left(Y_{u}+Y_{u}^{T}\right)+\left(\frac{3}{2}Y_{4}-Y_{4}^{T}\right)\epsilon$ in order to eliminate proton decay through the color triplet mediation. Note, however, that we cannot simultaneously suppress both sets of the scalar leptoquark couplings. This distinguishes our proposal from an approach advocated in Ref. Dvali:1992hc that insists on complete decoupling of the color triplet from the SM fermions, thus resulting, as specified in Ref. Dvali:1995hp , in existence of long-lived colored state(s).
  • If the color triplet $T\in 5_{H}$ is indeed light, it would form a complete $SU(5)$ multiplet with the SM Higgs boson doublet $D\in 5_{H}$ as far as the analysis of the running of the gauge coupling constants is concerned. Since the SM Higgs boson doublet is always beneficial for unification and aids the largeness of $M_{\mathrm{GUT}}$, the lightness of $T$ would consequentially lower maximal possible value of $M_{\mathrm{GUT}}$, thus making the proton decay signatures through the gauge boson mediation more accessible.
  • Suppression of the scalar mediation of proton decay signatures somewhat constrains Yukawa couplings of leptoquark $T$ to the SM fermions. (See, for example, Eqs. (20), (21), (22), and (23).) If one, furthermore, suppresses proton decay through gauge mediation, one would also introduce strong correlation between pairs of unitary matrices ($U$, $U_{c}$), ($D$, $E_{c}$), and ($E$, $D_{c}$). Consequently, the detection of the scalar leptoquark $T$ at current or future colliders would enable the verification of consistency between lepton-quark-leptoquark or quark-quark-leptoquark interactions and the observed final states.
  • Our approach to the ($D$)oublet-($T$)riplet splitting problem can be easily applied within the supersymmetric framework. For example, within the minimal supersymmetric $SU(5)$ model, augmented with higher-dimensional operators to address observed charged fermion masses, one of the two triplets would have couplings given in Eqs. (12), (13), and (14), whereas the other triplet would have couplings given in Eqs. (15) and (16). One could then suppress either the quark-quark or lepton-quark interactions of both of these triplets using the conditions already discussed in our manuscript. This, in turn, would completely suppress all $d=5$ and $d=6$ proton decay operators associated with these two triplets. We defer a full viability study of cancellation between $d=4$ and $d=5$ contributions towards charged fermion masses that is required if one is to implement our approach towards the ($D$)oublet-($T$)riplet splitting problem within the minimal supersymmetric $SU(5)$ model with higher-dimensional operators to future publication. (For several attempts to address the ($D$)oublet-($T$)riplet splitting problem within the minimal supersymmetric $SU(5)$ setting see Refs. Berezhiani:1998hg ; Bajc:2002bv ; Bajc:2002pg .)
VI Conclusions
We demonstrate that without additional assumptions about the flavor structure of the theory, the experimental data on partial proton decay lifetimes cannot meaningfully constrain the mass of the scalar leptoquark if one includes effects of higher-dimensional operators. Our result is applicable to any non-supersymmetric $SU(5)$ setup that $(i)$ uses a single $5_{H}$, where the aforementioned leptoquark resides, and $(ii)$ resorts to the use of $d=5$ operators to provide viable masses for the SM charged fermions. This undermines the underlying premise behind the need to implement the doublet-triplet mass splitting in this class of models.

### v4

@@ Abstract
We demonstrate that it is not necessary to substantially break the mass degeneracy between the Standard Model Higgs boson doublet and the corresponding scalar leptoquark, where these two fields comprise a single five-dimensional $SU(5)$ representation. More precisely, we show that the experimental data on partial proton decay lifetimes cannot place any meaningful bound on the mass of the scalar leptoquark, a.k.a. the color triplet, if one includes effects of higher-dimensional operators. We use the Georgi-Glashow model for our demonstration and point out that the complete suppression of proton decay via the scalar leptoquark mediation is not in conflict with the partial suppression of the gauge mediated two-body proton decay signatures. Finally, we derive a proper limit on the cutoff scale of the aforementioned higher-dimensional operators and outline how this scenario could be tested.
@@ I Introduction
One of the most prominent features of the the Georgi-Glashow model Georgi:1974sy is an ever-present need to introduce an enormous splitting between the masses of the two multiplets residing in the same five-dimensional scalar representation $5_{H}$ of the theory. One of these two multiplets is the Standard Model (SM) Higgs boson doublet $D$, whereas the other one is the color triplet $T$ that generates proton decay. Since the Higgs boson is observed to be light while the proton decay mediating scalar leptoquark should be extremely heavy, one needs to strongly break mass degeneracy between them. This issue has been referred to in the literature as the ($D$)oublet-($T$)riplet splitting problem Dimopoulos:1981zb ; Sakai:1981gr and it persists regardless of whether the theory is supersymmetric or not.
The crux of the ($D$)oublet-($T$)riplet splitting problem in non-supersymmetric $SU(5)$ is a need to precisely cancel, i.e., fine-tune, contributions to the Higgs boson mass from $d=4$ operators $m^{2}5^{i}_{H}5^{*}_{Hi}$, $m^{\prime}5^{i}_{H}24_{Hi}^{j}5^{*}_{Hj}$, $\lambda 5^{i}_{H}24_{Hi}^{j}24_{Hj}^{k}5^{*}_{Hk}$, and $\lambda^{\prime}5^{i}_{H}5^{*}_{Hi}24_{Hk}^{j}24_{Hj}^{k}$, where $i,j,k=1,\ldots,5$ are $SU(5)$ indices, $24_{H}$ is the scalar representation that breaks $SU(5)$ down to the SM gauge group Georgi:1974sy , and $m$, $m^{\prime}$, $\lambda$, and $\lambda^{\prime}$ are parameters of appropriate mass dimensions. The required cancellation takes place between contributions towards the mass of $D$ that are necessarily of the same order as the vacuum expectation value (VEV) of the SM singlet field behind the $SU(5)$ symmetry breaking in order for $T$ to remain extremely heavy.
There are several distinct approaches on how to overcome the issue of ($D$)oublet-($T$)riplet splitting within the four-dimensional $SU(5)$ framework. One approach is to implement a missing partner mechanism Masiero:1982fe whereby a triplet $T$ gets a large mass via coupling with an additional triplet field, whereas a doublet $D$ remains light due to an assumption that the common contributions towards masses of both $D$ and $T$ from terms such as $m^{2}5^{i}_{H}5^{*}_{Hi}$ are either all absent or very small. Note that this assumption is exactly opposite to the one that is required for the viability of the standard implementation of the ($D$)oublet-($T$)riplet splitting, but the end result is the same. There is a light multiplet $D$ and a heavy multiplet $T$, where both of these multiplets interact with the SM fermions. Since one needs to introduce additional triplet(s) without accompanying doublet(s), the missing partner mechanism implementation in $SU(5)$ calls for additional representations in the scalar sector of the theory to accomplish just that Masiero:1982fe .
Another possible approach to the ($D$)oublet-($T$)riplet splitting problem in $SU(5)$ is to decouple the triplet from the SM fermions, so that the triplet mass issue becomes irrelevant. Namely, triplet $T$ needs to be heavy if it yields proton decay. If one does not insist on unification of quarks and leptons of the SM, as in the original Georgi-Glashow proposal, one can couple triplet $T$ to the pairs of fermions, where one fermion is from the SM, whereas the other one is not. This can allow one to kinematically prevent proton decay in $SU(5)$, if these extra fermionic states are sufficiently heavy Fornal:2018aqc . This particular approach requires additional representations of both fermionic and scalar nature in order to be viable. The end result, in this case, is a triplet $T$ that can be light, but does not couple to the pairs of the SM fermions, or, at least, not to those that are relevant for preservation of matter stability.
@@
We will pursue an alternative direction that relies on inclusion of higher-dimensional operators and yields a triplet $T$ that exclusively couples to either the SM quark-quark or quark-lepton pairs but not to both. We will, for definiteness, use Georgi-Glashow model to provide a proof of existence of our observation although our analysis is also applicable to the $SU(5)$ theory that is broken with $75_{H}$ Hubsch:1984pg instead of $24_{H}$, the flipped $SU(5)$ scenario Barr:1981qv , and supersymmetric realisations of these unification models. We will explicitly demonstrate that the inclusion of $SU(5)$ invariant higher-dimensional operators allows one to completely suppress proton decay signatures via the scalar leptoquark mediation without affecting viability of the mechanism that generates masses of the SM charged fermions. We will furthermore show that this intriguing possibility is also compatible with the partial suppression of the gauge mediated two-body proton decay signatures.
@@
We introduce the most relevant details of the $SU(5)$ setup and the accompanying notation in Sec. II. The proton decay suppression mechanisms for both the scalar and gauge mediations are discussed in detail in Sec. III. We subsequently show how to establish an accurate upper bound on the cutoff scale associated with the higher-dimensional operators if one resorts to these suppression mechanisms in Sec. IV. Our findings are summarized in Sec. V.
@@ II Setup
$+\frac{1}{\Lambda}Y_{3}^{\alpha\beta}24_{Hm}^{n}\epsilon_{ijkln}+\frac{1}{\Lambda}Y_{4}^{\alpha\beta}24_{Hk}^{n}\epsilon_{ijlmn}\bigg{\}}+\mathrm{h.c.},$ (1)
@@
The VEVs of relevance are
@@
The SM fermion mass matrices are diagonalized via
@@ III Proton Decay Suppression
Since the linear combinations of Yukawa coupling matrices that enter the charged fermion masses $M_{E}$, $M_{D}$, and $M_{U}$, as given in Eqs. (5), (6), and (7), respectively, differ from those that are featured in the interactions of the scalar leptoquark $T$ with the SM fermions due to the presence of $d=5$ contributions, one can suppress the latter without affecting viability of the former. For example, one can set to zero all quark-quark couplings of $T$ in Eqs. (13) and (15) by imposing the following three relations
$Y_{2}\epsilon$ $=Y_{1}\epsilon-Y_{d},$ (17)
@@
where the viability of Eq. (18) will be discussed in detail later on. These relations consequentially lead to
@@
Even though we still need to demonstrate that one can indeed cancel $d=4$ contributions with $d=5$ contributions in the up-type quark sector with perturbative Yukawa couplings, Eqs. (17) through (23) already imply that generation of viable charged fermion masses is compatible with complete suppression of scalar mediated proton decay signatures. Therefore, one of the main results of this manuscript is that experimental data on partial proton decay lifetimes cannot provide any meaningful constraint on the scalar leptoquark mass if one includes effects of higher-dimensional operators.
If one is to include $d>5$ operators involving $5_{H}$ in the flavor sector of the theory, one would reintroduce quark-quark couplings with multiplet $T$. Note, however, that the viability of charged fermion masses as well as proton decay suppression can already be accommodated with $d=4$ and $d=5$ operators of Eq. (1). This simply means that the proton decay data can only provide constraint on the coefficients associated with $d>5$ operators as a function of scalar leptoquark mass without invalidating our result. Also, one can simply redefine interactions obtained at the $d=5$ level by incorporating the couplings generated though $d>5$ operators and proceed with aforementioned cancellations. In other words, the conditions in Eqs. (17), (18), and (19) represent one possible flavor scenario, among many, that does not yield proton decay through the scalar exchange and thus serves as a proof of existence we are interested in. Our proposal relies on precise cancellation, i.e., fine-tuning, as the standard implementation of the ($D$)oublet-($T$)riplet splitting but, in our case, the cancellation takes place in the flavor sector instead of the scalar sector and yields potentially light triplet $T$ that still couples to the SM fermions.
We have explicitly demonstrated how to suppress the quark-quark couplings of the scalar $T$ through cancellations of individual contributions towards its interactions with the SM fermions. We could have alternatively opted to suppress the quark-lepton couplings of Eqs. (12), (14), and (16) via the following two relations
$Y_{2}\epsilon$ $=\frac{2}{3}\left(Y_{d}-Y_{1}\epsilon\right),$ (24)
$\left(Y_{3}+Y_{3}^{T}\right)\epsilon$ $=\left(Y_{u}+Y_{u}^{T}\right)+\left(\frac{3}{2}Y_{4}-Y_{4}^{T}\right)\epsilon.$ (25)
This is yet another way to eliminate proton decay through the color triplet mediation. Note, however, that we cannot simultaneously suppress both sets of the scalar leptoquark couplings. (This, for example, distinguishes our proposal from an approach advocated in Ref. Dvali:1992hc and implemented within $SO(10)$ framework that insists on complete decoupling of the color triplet from the SM fermions, thus resulting in possible existence of long-lived colored state(s) Dvali:1995hp .)
Our approach to the ($D$)oublet-($T$)riplet splitting problem, as we already stated, can be easily implemented within the supersymmetric framework. For example, within the minimal supersymmetric $SU(5)$ model, augmented with higher-dimensional operators to address observed charged fermion masses, one of the two triplets would have couplings given in Eqs. (12), (13), and (14), whereas the other triplet would have couplings given in Eqs. (15) and (16). One could then suppress either the quark-quark or lepton-quark interactions of both of these triplets using the conditions already discussed in our manuscript. This, in turn, would completely suppress all $d=5$ and $d=6$ proton decay operators associated with these two triplets. (There have been several attempts Berezhiani:1998hg ; Bajc:2002bv ; Bajc:2002pg within the minimal supersymmetric $SU(5)$ to exploit the freedom in the flavor sector of the theory to try to partially suppress couplings of the triplets to the SM fermions in order to alleviate tension between experimentally established proton decay lifetime limits and associated theory predictions.)
@@
Although proton decay via gauge boson exchange cannot be completely rotated away in $SU(5)$ Nandi:1982ew , it is possible to significantly suppress coefficients in Eqs. (29), (30), and (31). This, for example, can be achieved with the followings set of conditions Dorsner:2004xa
@@
It is clear that simultaneous suppression of both the scalar and gauge mediation of proton decay does not only constrain Yukawa couplings of leptoquark $T$ to the SM fermions, but also introduces strong correlation between pairs of unitary matrices ($U$, $U_{c}$), ($D$, $E_{c}$), and ($E$, $D_{c}$). The detection of the scalar leptoquark $T$ decay patterns at current or future colliders would consequently enable verification of consistency between predicted lepton-quark-leptoquark or quark-quark-leptoquark interactions and the observed final states.
@@ IV Cutoff Scale Limit
For our purposes it is sufficient to take $M_{U}^{\mathrm{diag}}=\mathrm{diag}(0,0,m_{t})=\mathrm{diag}(0,0,y_{t}v_{5}/\sqrt{2})$, where $m_{t}$ and $y_{t}$ are the top quark mass and Yukawa coupling at $M_{\mathrm{GUT}}$ scale, respectively. (Our derivation of the limit on $\Lambda$ is accurate up to the corrections of the order of $m_{c}/m_{t}$, where $m_{c}$ is the charm quark mass.) This decomposition allows us to find a lower limit on the largest possible entry in $A$, i.e., $\max|(A)_{\alpha\beta}|$, in units of $y_{t}$. Since the elements of $A$ can only originate from the very last operator in Eq. (7), our approach should yield an accurate lower limit on parameter $\epsilon\equiv v_{24}/\Lambda$ if we insist on perturbativity of Yukawa couplings. Coincidentally, decomposition of Eq. (36) is compatible with the conditions in Eqs. (18) and (19). This, in turn, will enable us to explicitly demonstrate that one can cancel $d=4$ contributions with $d=5$ contributions with perturbative couplings in the up-type quark sector.
@@
Here, $c_{\theta}\equiv\cos\theta$ and $s_{\theta}\equiv\sin\theta$, where $\theta$ is a free parameter, while $\eta_{1}$, $\eta_{2}$, $\eta_{3}$, and $\eta_{4}$ are four arbitrary phases. (The values of elements in $V_{\mathrm{CKM}}$ are sourced from the Particle Data Group ParticleDataGroup:2022pth , although the only $V_{\mathrm{CKM}}$ entry that matters for our analysis is the Cabibbo angle.) We can, furthermore, represent unitary matrix $U$ via
@@
To proceed, we need to establish what the viable values of $v_{24}$ and $y_{t}$, to be used as input in Eq. (41), are. To do that we first observe that $v_{24}=M_{\mathrm{GUT}}\sqrt{(2\alpha_{\mathrm{GUT}}^{-1})/25\pi}$, where $\alpha_{\mathrm{GUT}}=g_{\mathrm{GUT}}^{2}/(4\pi)$. What we thus need are values of gauge coupling unification scale $M_{\mathrm{GUT}}$ and associated values for $\alpha_{\mathrm{GUT}}$ and $y_{t}$.
To obtain appropriate values of $M_{\mathrm{GUT}}$, $\alpha_{\mathrm{GUT}}$, and $y_{t}$ we resort to a model Dorsner:2005fq ; Dorsner:2005ii ; Dorsner:2006hw that $(i)$ uses $d=4$ operator to generate neutrino masses with a single fifteen-dimensional scalar representation through the type II seesaw mechanism Lazarides:1980nt ; Mohapatra:1980yp , $(ii)$ relies on operators of Eq. (1) to generate charged fermion masses, and $(iii)$ implements Eqs. (32), (33), and (34) to suppress gauge mediated proton decay signatures. This model yields $M_{\mathrm{GUT}}=4.5\times 10^{14}$ GeV, $\alpha_{\mathrm{GUT}}^{-1}=37.1$ and $y_{t}=0.42$, where $\alpha_{\mathrm{GUT}}$ is evaluated via the two-loop level gauge coupling unification analysis, whereas $y_{t}$, at $M_{\mathrm{GUT}}$, is obtained through the one-loop level renormalisation group equation running. If these values for $\alpha_{\mathrm{GUT}}$ and $y_{t}$ are inserted into Eq. (41) we find that the upper limit on $\Lambda$ for a model of Ref. Dorsner:2005fq reads $\Lambda<57M_{\mathrm{GUT}}$, where we assume that $5/2\max|(Y_{4})_{\alpha\beta}|=\sqrt{4\pi}$. This bound on $\Lambda$, as inferred from Eq. (41), is another key finding of this manuscript. It is, as already advocated, much more stringent than the $\Lambda\lesssim 900M_{\mathrm{GUT}}$ bound reported in Ref. Dorsner:2006hw .
We note that it has been recently argued Senjanovic:2024uzn that the particle content of the Georgi-Glashow model Georgi:1974sy is already sufficient to address gauge coupling unification if one adds $d=5$ operators in the gauge sector of the theory Shafi:1983gz . That proposal Senjanovic:2024uzn also relies on suppression of gauge mediated proton decay signatures for its viability but uses $d=5$ operator to accommodate neutrino masses. We find that the most stringent cutoff scale limit of that model actually originates from the need to have viable mass scale for neutrinos. Consequentially, the use of Eq. (41) to find associated cutoff within that model would not be appropriate.
@@
Let us, as a qualitative example, assume that $\epsilon=0.05$. This, for $\alpha_{\mathrm{GUT}}^{-1}=37.1$ and $y_{t}=0.42$, implies that $\Lambda=19.4M_{\mathrm{GUT}}=8.7\times 10^{15}$ GeV, $2.25\max|(Y_{3}+Y_{3}^{T})_{\alpha\beta}|>$ 2.2, $2.45\max|(Y_{4})_{\alpha\beta}|/2>$ 2.4, and $0.882\max|(Y_{u}+Y_{u}^{T})_{\alpha\beta}|>$ 0.88 if one is to implement suppression of both scalar and gauge mediations of proton decay within a model of Refs. Dorsner:2005fq ; Dorsner:2005ii ; Dorsner:2006hw .
@@
V Conclusions
We demonstrate that without additional assumptions about the flavor structure of the theory, the experimental data on partial proton decay lifetimes cannot meaningfully constrain the mass of the scalar leptoquark in $SU(5)$ if one includes effects of higher-dimensional operators. Our result is applicable to both supersymmetric and non-supersymmetric $SU(5)$ frameworks that resort to the use of $d=5$ operators to provide viable masses for the SM charged fermions. This undermines the underlying premise behind the need to implement the doublet-triplet mass splitting in this class of models.